# FLINT: Learning-based Flow Estimation and Temporal Interpolation for Scientific Ensemble Visualization

### v1

@@ I Introduction
In summary, we propose FLINT, a new flexible method for the estimation of flow for 2D+time and 3D+time scientific ensembles. The FLINT code will be available upon acceptance of the publication. We consider our main contributions to be as follows:
  1. {old}
    To the best of our knowledge, FLINT is the first approach that not only generates high-quality temporal interpolants between scalar fields (like density or luminance), but also estimates a corresponding flow field for each intermediate. To the best of our knowledge, FLINT is the first approach to perform flow estimation for scientific ensembles, estimating a corresponding flow field for each discrete time point, even in the absence of original flow information.
@@
  3. FLINT can handle both 2D+time and 3D+time ensembles, and does not require domain-specific assumptions, complex pre-training, or intermediate fine-tuning on simplified datasets.
  4. FLINT is capable of high-quality flow estimation in cases (1) in which flow information is partially available (flow-supervised, e.g., simulations where the flow is omitted due to space constraints) or (2) not available at all (flow-unsupervised, e.g., in experimental image sequences captured via cameras). {old}
@@ II Related Work
Flow field estimation in scientific visualization. Kappe et al. [13] reconstructed and visualized 3D local flow from 3D+time microscopy data using both image processing and optical flow methods. Kumpf et al. [14] utilized ensemble sensitivity analysis (ESA) to obtain the evolution of sensitivity features in 2D or 3D geo-spatial data. They used forward and backward optical flow-based feature assignment for tracing the evolution of sensitivities through time. Manandhar et al. [15] developed a 3D optical flow estimation method for light microscopy image volumes. A dense 3D flow field was obtained from the 3D displacement vectors in a pair of such volumes. However, these approaches involve case-specific assumptions or rely on ground truth (GT) from synthetic data, limiting their generalizability to scientific ensembles derived from physical experiments. In contrast, FLINT does neither require data-specific assumptions nor available GT flow field.
@@
Machine learning-based upscaling & super-resolution. ML-based upscaling and super-resolution approaches have gained significant attention across various domains, including image processing, computer vision, and scientific visualization [25, 26]. They can be categorized into three main groups: spatial super-resolution (SSR), temporal super-resolution (TSR), and spatio-temporal super-resolution (STSR). SRCNN [27], SRFBN [28], and SwinIR [29] are SSR techniques that aim to improve the spatial resolution of images by generating realistic and fine-grained details. TSR techniques focus on interpolating intermediate timesteps from subsampled time sequences while maintaining the spatial resolution. Methods such as phase-based interpolation [30], SepConv [31], and SloMo [32] tackle the challenge of synthesizing high-quality intermediate frames to enhance the temporal resolution of videos. While previous approaches, like STNet [33], can upscale data with a fixed spatial or temporal scale factor, they do not simultaneously address both spatial and temporal super-resolution. Han et al. proposed a recurrent generative network, TSR-TVD [34], designed to be trained on one variable and then applied to generate a temporally higher-resolution version of another variable from its lower-resolution counterpart. Recent methods like Filling the Void [35], SSR-TVD [36], FFEINR [37], HyperINR [38], CoordNet [39], and STSR-INR [40] have made significant strides in enabling TSR or SSR of data fields at arbitrary resolution. While FLINT achieves state-of-the-art accuracy in temporal interpolation between two given fields, its primary objective and notable strength lies in flow estimation. Our method not only excels in reconstructing temporal data with high precision but also produces impressive results in supplementing the data with accurate flow information.
@@ III Method
FLINT implements a student-teacher architecture [44] which has several key advantages. It enables more accurate predictions as it is trained on GT temporal scalar fields (as demonstrated in Sec. V). Guidance through the teacher model further enhances the student model’s learning process, resulting in stable training, faster convergence, and improved model robustness. FLINT does not require pre-training and intermediate fine-tuning on simplified datasets; it already converges on the target datasets, in contrast to previous comparable works in computer vision [5, 6, 7]. FLINT further yields significantly faster convergence and achieves an accuracy of more than 90% within the initial 30% of the total training time. Below, we describe the neural network architecture employed in FLINT (Sec. III-A), and present the temporal interpolation and flow estimation pipeline, both for the training and inference phases (Sec. III-B). Our loss function for training can flexibly adapt to different scenarios as described in Sec. III-C. FLINT builds upon RIFE [8] and modifies as well as extends it in several ways; a detailed comparison is presented in Sec. III-D.
@@ III-A FLINT Network Architecture
The network architecture is a feed-forward CNN, see Fig. 2. The student network consists of $N$ stacked convolutional blocks (Conv Block), each incorporating convolutional (Conv) and deconvolutional (Deconv) layers (the middle column in the orange box of Fig. 2 shows an expanded view of a convolution block). There are 256 feature channels in all convolutional layers of the first block, 192 of the second and third, and 128 of the last block. The number of channels in the layers of the teacher block Conv Blockteach is set to 128, similarly to the last block of the student model. We utilize a PReLU activation function [45] in all layers except for the last one. The teacher model crucially features a dedicated Conv Blockteach which enables it to directly consider a GT scalar field by receiving $D^{\mbox{{{\it{GT}}}}}_{t}$ through an additional channel. Note that traditionally such information is not available directly to the network architecture but is only considered in the loss function. The loss components that drive the training process are shown on the right side in Fig. 2. Student and teacher models are jointly optimized during training (i.e., following an online scheme [44]), with the teacher model refining the student model’s results. The proposed FLINT model architecture serves as the foundation for tasks with and without available GT flow fields, differing only in the applied loss functions (Sec. III-C).
@@
We employ the student-teacher model approach as it presents several key advantages over relying solely on the reconstruction loss calculated from the available $D^{\mbox{{{\it{GT}}}}}_{t}$. It enables more accurate predictions as it is trained on ground-truth {old}timestep temporal fields (as demonstrated in Sec. V). Guidance through the teacher model further enhances the student model’s learning process, resulting in stable training, faster convergence, and improved model robustness. FLINT does not require pre-training and intermediate fine-tuning on simplified datasets, as it achieves convergence already on the target datasets, in contrast to previous comparable works in computer vision [5, 6, 7]. Notably, compared to these methods, FLINT demonstrates a significantly shorter convergence time, also achieving an accuracy of more than 90% within the initial 30% of the total training time.
@@ III-B Flow Estimation and Scalar Field Interpolation
Refining the intermediate flows. The intermediate flow fields are computed via $N$ convolutional blocks by iterative refinement, see Fig. 2. Superscripts $i$ denote the various quantities at iteration $i$. Given two input fields $D_{s}$ and $D_{u}$ and a timepoint $t$, with $s<t<u$, Conv Block0 computes a rough prediction of intermediate flow fields ${\hat{F}^{0}_{t\rightarrow s}},{\hat{F}^{0}_{t\rightarrow u}}$ and fusion mask $M^{0}$ to capture large motions. Then in Conv Blocki, $i>0$, $D_{s}$ and $D_{u}$ are first backward-warped using Eq. (1), based on the intermediate flows ${\hat{F}^{i-1}_{t\rightarrow s}},{\hat{F}^{i-1}_{t\rightarrow u}}$ and mask $M^{i-1}$ of the previous iteration (see the yellow Warping block in Fig. 2, middle column). Next, $D_{s}$ and $D_{u}$, warped frames ${\hat{D}^{i-1}_{t\leftarrow s}}$ and ${\hat{D}^{i-1}_{t\leftarrow u}}$, intermediate flows ${\hat{F}^{i-1}_{t\rightarrow s}}$ and ${\hat{F}^{i-1}_{t\rightarrow u}}$, mask $M^{i-1}$, and timestep $t$ are concatenated and processed by the stack of convolution and deconvolution layers in Conv Blocki. This results in updated ${\hat{F}^{i}_{t\rightarrow s}},{\hat{F}^{i}_{t\rightarrow u}}$, and $M^{i}$, which then enter the next Conv Block. The process continues until the last Conv BlockN-1 has finished computation, producing final estimates ${\hat{F}^{N-1}_{t\rightarrow s}},{\hat{F}^{N-1}_{t\rightarrow u}}$, and $M^{N-1}$.
@@ III-C Loss Function
where $F^{\mbox{{{\it{GT}}}}}_{t}$ is the GT flow at time $t$, $\hat{F}^{i}_{t}$ is the flow output from the corresponding $i^{th}$ block of the student network (Eq. (2b)), $N=4$ is the number of blocks in the model, and $\hat{F}_{t}^{teach}$ is the flow output from the teacher block. We experimentally established the value of $\gamma$ as 0.8, aligning with the RAFT loss and validating this choice through our hyperparameter search. Again, $F^{\mbox{{{\it{GT}}}}}_{t}$ is only used for calculation of the error; it is never used as input to the student-teacher network.
@@
$\mathcal{L}_{dis}=\sum_{j\in\{s,u\}}\lVert\color[rgb]{0.,0.,0.}\definecolor[named]{pgfstrokecolor}{rgb}{0.,0.,0.}\hat{F}^{N-1}_{t\rightarrow j}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\hat{F}_{t\rightarrow j}^{teach}\rVert_{2},$ (7)
@@
The photometric loss in our experiments is the error between a field of the first timestep and the backward warping of a field of the second timestep, which was interpolated based on the estimated flow. We also conducted experiments incorporating the smoothness loss component, commonly used in optical flow estimation tasks [24]. However, in our case, where there are generally no clearly distinguishable objects within our ensembles, the smoothness loss did not improve the results. The metric results obtained by incorporating the smoothness loss component are reported in Table V under the “FLINT smooth” row. The Charbonnier penalty function is used in both photometric and smoothness loss components to provide robustness against outliers [46]. The photometric loss is computed as the difference between the fields $D_{s},D_{u}$, and the {old}backward/inverse warped field of the reconstructed field $\hat{D}_{t}$:
@@ III-D Comparison between FLINT and RIFE methods
Similarities. RIFE and FLINT both receive two scalar input fields and employ several convolution blocks to produce approximate intermediate flows and a fusion mask. Warped fields are computed from the input frames with help of the intermediate flows. Merging the two warped frames with help of the fusion mask via Eq. (2a) yields the interpolated field. RIFE and FLINT feature an online end-to-end trainable student-teacher architecture using a special teacher convolution block. To calculate the model residual error, both use an optimization process that minimizes a loss function, consisting of various terms.
Improvements and Extensions. In RIFE the number of student convolution blocks is fixed to three, while in FLINT we can flexibly adapt the number and configuration of blocks for the most optimal performance. Furthermore, RIFE’s use of different scales for different blocks with bilinear resize is replaced by convolution and deconvolutional layers, resulting in more learnable parameters. Additionally, FLINT comprises convolutional and deconvolutional layers with varying strides in each block, unlike RIFE’s eight convolutional layers with a stride of one per block. This autoencoder-like design effectively captures key input features [48], as demonstrated by the performance gains shown in Table V.
@@
Dens GT Dens FLINT Dens diff ×$\times 100$ Dens RIFE Dens diff ×$\times 100$ Flow GT Flow FLINT Flow diff Flow RIFE Flow diff RIFE Stream GT Stream FLINT Refer to caption
@@
Dens GT Dens FLINT Dens diff ×$\times 100$ Dens RIFE Dens diff ×$\times 100$ Flow GT Flow FLINT Flow diff Flow RIFE Flow diff RIFE Stream GT Stream FLINT Refer to caption
@@
Figure 4: LBS ensemble: FLINT flow field estimation and temporal density interpolation at the timesteps with an asterisk (*)—(a) 6$\times$ and (b) 18$\times$ interpolation. Each subfigure, from top to bottom shows GT density, FLINT interpolated density, density difference (magnified by ×$\times 100$); GT flow, FLINT flow prediction, difference with GT flow, flow estimated by RIFE, difference with GT flow, GT streamlines, and FLINT streamlines. The colorbar on the top right is used for mapping density, and the one on the bottom right maps flow magnitude.

### v2

@@ I Introduction
In summary, we propose FLINT, a new flexible method for the estimation of flow for 2D+time and 3D+time scientific ensembles. The FLINT code will be made available publicly. We consider our main contributions to be as follows:
  1. To the best of our knowledge, FLINT is the first approach to achieve high-quality flow estimation in scenarios (1) where flow information is partially available (flow-supervised, e.g., simulations where flow is omitted due to storage constraints) or (2) entirely unavailable (flow-unsupervised, e.g., experimental image sequences captured via cameras).
@@
  3. FLINT can handle both 2D+time and 3D+time ensembles without requiring domain-specific assumptions, complex pre-training, or intermediate fine-tuning on simplified datasets.
    {old}
@@ II Related Work
Flow field estimation in scientific visualization. Kappe et al. [13] reconstructed and visualized 3D local flow from 3D+time microscopy data using both image processing and optical flow methods. Kumpf et al. [14] utilized ensemble sensitivity analysis (ESA) to obtain the evolution of sensitivity features in 2D or 3D geo-spatial data. They used forward and backward optical flow-based feature assignment for tracing the evolution of sensitivities through time. Manandhar et al. [15] developed a 3D optical flow estimation method for light microscopy image volumes. A dense 3D flow field was obtained from the 3D displacement vectors in a pair of such volumes. Gu et al. introduced Scalar2Vec [16]—a deep learning-based framework that translates scalar fields into velocity vector fields, using a k-complete bipartite translation network (kCBT-Net). We conceptually compare our findings to Scalar2Vec in Sec. VI. However, these approaches involve case-specific assumptions [13, 14, 15] or rely on ground truth (GT) [16] from synthetic data, limiting their generalizability to scientific ensembles derived from physical experiments. In contrast, FLINT does neither require data-specific assumptions nor an available GT flow field.
@@
Machine learning-based upscaling & super-resolution. ML-based upscaling and super-resolution approaches have gained significant attention across various domains, including image processing, computer vision, and scientific visualization [26, 27]. They can be categorized into three main groups: spatial super-resolution (SSR), temporal super-resolution (TSR), and spatio-temporal super-resolution (STSR). SRCNN [28], SRFBN [29], and SwinIR [30] are SSR techniques that aim to improve the spatial resolution of images by generating realistic and fine-grained details. TSR techniques focus on interpolating intermediate timesteps from subsampled time sequences while maintaining the spatial resolution. Methods such as phase-based interpolation [31], SepConv [32], and SloMo [33] tackle the challenge of synthesizing high-quality intermediate frames to enhance the temporal resolution of videos. While previous approaches, like STNet [34], can upscale data with a fixed spatial or temporal scale factor, they do not simultaneously address both spatial and temporal super-resolution. Han et al. proposed a recurrent generative network, TSR-TVD [35], designed to be trained on one variable and then applied to generate a temporally higher-resolution version of another variable from its lower-resolution counterpart. Recent methods like Filling the Void [36], SSR-TVD [37], FFEINR [38], HyperINR [39], CoordNet [40], and STSR-INR [41] have made significant strides in enabling TSR or SSR of data fields at arbitrary resolution. While FLINT achieves state-of-the-art accuracy in temporal interpolation between two given fields, its primary objective and notable strength lies in flow estimation. This task involves predicting flow fields directly from given scalar fields, assuming no flow information is available during the inference stage. Methods like CoordNet or STSR-INR, which are powerful in TSR or SSR tasks, are not designed for predicting vector fields from scalar ones. Our FLINT method not only excels in reconstructing temporal data with high precision but also produces impressive results in supplementing the data with accurate flow information.
@@ III Method
FLINT implements a student-teacher architecture [45] which has several key advantages. It enables more accurate estimations as it is trained on GT temporal scalar fields (as demonstrated in Sec. V). Guidance through the teacher model further enhances the student model’s learning process, resulting in stable training, faster convergence, and improved model robustness. FLINT does not require pre-training and intermediate fine-tuning on simplified datasets; it already converges on the target datasets, in contrast to previous comparable works in computer vision [5, 6, 7]. FLINT further yields significantly faster convergence and achieves an accuracy of more than 90% within the initial 30% of the total training time. Below, we describe the neural network architecture employed in FLINT (Sec. III-A), and present the temporal interpolation and flow estimation pipeline, both for the training and inference phases (Sec. III-B). Our loss function for training can flexibly adapt to different scenarios as described in Sec. III-C. FLINT builds upon RIFE [8] and modifies as well as extends it in several ways; a detailed comparison is presented in Sec. III-D.
@@ III-A FLINT Network Architecture
The network architecture is a feed-forward CNN, see Fig. 2. The student network consists of $N$ stacked convolutional blocks (Conv Block), each incorporating convolutional (Conv) and deconvolutional (Deconv) layers (the middle column in the orange box of Fig. 2 shows an expanded view of a convolution block). There are 256 feature channels in all convolutional layers of the first block, 192 of the second and third, and 128 of the last block. The number of channels in the layers of the teacher block Conv Blockteach is set to 128, similarly to the last block of the student model. We utilize a PReLU activation function [46] in all layers except for the last one. The teacher model crucially features a dedicated Conv Blockteach which enables it to directly consider a GT scalar field by receiving $D^{\mbox{{{\it{GT}}}}}_{t}$ through an additional channel. Note that traditionally such information is not available directly to the network architecture but is only considered in the loss function. The loss components that drive the training process are shown on the right side in Fig. 2. Student and teacher models are jointly optimized during training (i.e., following an online scheme [45]), with the teacher model refining the student model’s results. According to our experiments, this streamlined one-stage online training approach significantly reduces training time compared to two-stage optimization, where the teacher model is trained first, and a student network is subsequently trained to align with the teacher’s outputs. As demonstrated in Sec. VI-C, our one-stage approach does not yield any degradation in performance while significantly reducing training time. The proposed FLINT model architecture serves as the foundation for tasks with and without available GT flow fields, differing only in the applied loss functions (Sec. III-C).
@@
We employ the student-teacher model approach as it presents several key advantages over relying solely on the reconstruction loss calculated from the available $D^{\mbox{{{\it{GT}}}}}_{t}$. It enables more accurate estimations as it is trained on ground-truth {old}timestep temporal fields (as demonstrated in Sec. V). Guidance through the teacher model further enhances the student model’s learning process, resulting in stable training, faster convergence, and improved model robustness. FLINT does not require pre-training and intermediate fine-tuning on simplified datasets, as it achieves convergence already on the target datasets, in contrast to previous comparable works in computer vision [5, 6, 7]. Notably, compared to these methods, FLINT demonstrates a significantly shorter convergence time, also achieving an accuracy of more than 90% within the initial 30% of the total training time.
@@ III-B Flow Estimation and Scalar Field Interpolation
Refining the intermediate flows. The intermediate flow fields are computed via $N$ convolutional blocks by iterative refinement, see Fig. 2. This coarse-to-fine process iteratively and jointly refines both the flow fields and the fusion mask at each Conv Block, ensuring progressively consistent and high-quality updates throughout the network. Superscripts $i$ in Fig. 2 denote the various quantities at iteration $i$. Given two input fields $D_{s}$ and $D_{u}$ and a timepoint $t$, with $s<t<u$, Conv Block0 computes a rough estimation of intermediate flow fields ${\hat{F}^{0}_{t\rightarrow s}},{\hat{F}^{0}_{t\rightarrow u}}$ and fusion mask $M^{0}$ to capture large motions. Then in Conv Blocki, $i>0$, $D_{s}$ and $D_{u}$ are first backward-warped using Eq. (1), based on the intermediate flows ${\hat{F}^{i-1}_{t\rightarrow s}},{\hat{F}^{i-1}_{t\rightarrow u}}$ and mask $M^{i-1}$ of the previous iteration (see the yellow Warping block in Fig. 2, middle column). Next, $D_{s}$ and $D_{u}$, warped frames ${\hat{D}^{i-1}_{t\leftarrow s}}$ and ${\hat{D}^{i-1}_{t\leftarrow u}}$, intermediate flows ${\hat{F}^{i-1}_{t\rightarrow s}}$ and ${\hat{F}^{i-1}_{t\rightarrow u}}$, mask $M^{i-1}$, and timestep $t$ are concatenated and processed by the stack of convolution and deconvolution layers in Conv Blocki. This results in updated ${\hat{F}^{i}_{t\rightarrow s}},{\hat{F}^{i}_{t\rightarrow u}}$, and $M^{i}$, which then enter the next Conv Block. The process continues until the last Conv BlockN-1 has finished computation, producing final estimates ${\hat{F}^{N-1}_{t\rightarrow s}},{\hat{F}^{N-1}_{t\rightarrow u}}$, and $M^{N-1}$.
@@ III-C Loss Function
where $F^{\mbox{{{\it{GT}}}}}_{t}$ is the GT flow at time $t$, $\hat{F}^{i}_{t}$ is the flow output from the corresponding $i^{th}$ block of the student network (Eq. (2b)), $N=4$ is the number of blocks in the model, and $\hat{F}_{t}^{teach}$ is the flow output from the teacher block. We experimentally established the value of $\gamma$ as 0.8, aligning with the RAFT loss and validating this choice through our hyperparameter search.
Again, $F^{\mbox{{{\it{GT}}}}}_{t}$ is only used for calculation of the error; it is never used as input to the student-teacher network.
@@
$\mathcal{L}_{dis}=\sum_{j\in\{s,u\}}\lVert\hat{F}^{N-1}_{t\rightarrow j}-\hat{F}_{t\rightarrow j}^{teach}\rVert_{2},$ (7)
@@
The photometric loss in our experiments is the error between a field of the first timestep and the backward warping of a field of the second timestep, which was interpolated based on the estimated flow. We also conducted experiments incorporating the smoothness loss component, commonly used in optical flow estimation tasks [25]. However, in our case, where there are generally no clearly distinguishable objects within our ensembles, the smoothness loss did not improve the results. The metric results obtained by incorporating the smoothness loss component are reported in Table VI under the “FLINT smooth” row. The Charbonnier penalty function is used in both photometric and smoothness loss components to provide robustness against outliers [47]. The photometric loss is computed as the difference between the fields $D_{s},D_{u}$, and the {old}backward/inverse warped field of the reconstructed field $\hat{D}_{t}$:
@@ III-D Comparison between FLINT and RIFE methods
Similarities. RIFE and FLINT both receive two scalar input fields and employ several convolution blocks to produce approximate intermediate flows and a fusion mask. Warped fields are computed from the input frames with the help of the intermediate flows. Merging the two warped frames with the help of the fusion mask via Eq. (2a) yields the interpolated field. RIFE and FLINT feature an online end-to-end trainable student-teacher architecture using a special teacher convolution block. To calculate the model residual error, both use an optimization that minimizes a loss function, consisting of various terms.
Improvements and Extensions. In RIFE the number of student convolution blocks is fixed to three, while in FLINT we can flexibly adapt the number and configuration of blocks for the most optimal performance. Furthermore, RIFE’s use of different scales for different blocks with bilinear resize is replaced by convolution and deconvolutional layers, resulting in more learnable parameters. Additionally, FLINT comprises convolutional and deconvolutional layers with varying strides in each block, unlike RIFE’s eight convolutional layers with a stride of one per block. This autoencoder-like design effectively captures key input features [49], as demonstrated by the performance gains shown in Table VI.
@@
IV Study Setup
In this section, we describe the training setup, provide an overview of the datasets used in our experiments and discuss the evaluation methods employed to assess the FLINT results.
IV-A Training
We apply the standard prepossessing step of normalization to [0, 1] when necessary before the start of the training. FLINT is optimized using AdamW [51] — an adaptive gradient descent method with weight decay used in back-propagation algorithms for training feed-forward neural networks that combines the benefits from both the Adam optimizer [52] and $L_{2}$-regularization. We employ early stopping with a patience parameter of 30, which is equivalent to regularization [53] and helps to prevent overfitting on the training data. We use an experimentally determined learning rate of $6\times 10^{-4}$ for the 2D case and $1\times 10^{-4}$ for the 3D case respectively with a cosine annealing scheduler that gradually decreases the learning rate to $6\times 10^{-6}$ and $1\times 10^{-6}$ respectively by the end of the training. We train FLINT with mini-batches of size 32 for both {old}the LBS and Droplets ensembles 2D ensemble datasets and mini-batches of size 2 for 3D datasets used in our study (see Sec. IV-B). We split the set of all available data into training, validation (for monitoring training progress), and test subsets.
Note that training and test data are obtained by subsampling the original dataset, which is assumed to provide the GT scalar fields. Then FLINT performs interpolation at missing timesteps.
IV-B Datasets
We consider four scientific datasets in our study.
Nyx (flow-supervised). The fourth dataset is a 3D+time ensemble based on the compressible cosmological hydrodynamics simulation Nyx, developed by Lawrence Berkeley National Laboratory [55]. We consider an ensemble comprising 18 members, each consisting of a maximum of 1600 timesteps with a spatial resolution of $128\times 128\times 128$. It contains density and the $x$, $y$, $z$ components of velocity. Akin to InSituNet [56], we vary three parameters for ensemble generation: the total matter density ($\Omega_{m}\in[0.12,0.155]$), the total density of baryons ($\Omega_{b}\in[0.0215,0.0235]$), and the Hubble constant ($h\in[0.55,0.7]$). We randomly sample a training subset of 500 timesteps after selecting every 5th timestep across different ensemble members for training, validation, and testing.
To visualize a simulation ensemble field, we use the Turbo colormap [57]. We visualize the 2D flow field via arrow glyphs and 3D flow via volume rendering with a transfer function that indicates flow direction. Examples of the LBS and Droplets ensemble fields, as well as 5Jets and Nyx transfer functions, are in the supplementary material.
Dens GT Dens FLINT Dens diff ×$\times 100$ Dens RIFE Dens diff ×$\times 100$ Flow GT Flow FLINT Flow diff Flow RIFE Flow diff RIFE Pathline GT Pathline FLINT Refer to caption
@@ IV-B Datasets
Dens GT Dens FLINT Dens diff ×$\times 100$ Dens RIFE Dens diff ×$\times 100$ Flow GT Flow FLINT Flow diff Flow RIFE Flow diff RIFE Pathline GT Pathline FLINT Refer to caption
@@
Figure 4: LBS ensemble: FLINT flow field estimation and temporal density interpolation at the timesteps with an asterisk (*)—(a) 6$\times$ and (b) 18$\times$ interpolation. Each subfigure, from top to bottom shows GT density, FLINT interpolated density, difference with GT density (magnified by ×$\times 100$), RIFE interpolated density, difference with GT density (magnified by ×$\times 100$); GT flow, FLINT flow estimation, difference with GT flow, flow estimated by RIFE, difference with GT flow, GT pathlines, and FLINT pathlines. The colorbar on the top right maps density, and the one on the bottom right maps flow magnitude.
IV-C Evaluation
{old}